OpenStudy (anonymous):

y=3x+3 2y= 2/3x +12 how to solve to systems of equation?

OpenStudy (scorcher219396):

Isolate one of the variables of one of the equations (doesn't matter what variable or equation, but some ways will be easier than others) Your y is already isolated, now can substitute it into the other equation. Wherever there is a y in the second equation, put in what is on the right of the first equation 2(3x+3)=2/3x +12 Then solve for x Once you get that x, plug it back into one of the equations (doesn't matter which, again some ways will be easier than others) and solve for y then

OpenStudy (anonymous):

6x+6 = 2/3x +12. I am confuse what to do next tho?

OpenStudy (anonymous):

Do I take the 6 out or 6x?

OpenStudy (dan815):

first put all the x on one side

OpenStudy (dan815):

and everything else on the other side

OpenStudy (anonymous):

Do I subract 2/3 to 6x

OpenStudy (dan815):

okay do that to both sides of the equation what happens?

OpenStudy (anonymous):

Wait do I need to find common bottom for this one?

OpenStudy (dan815):

yes

OpenStudy (anonymous):

I got 6/3.

OpenStudy (dan815):

\[6x+6\color \red {-6} = \frac{2x}{3} +12\color \red {-6} \\ 6x = \frac{2x}{3} +6\\ 6x-\color \red{\frac{2x}{3}} = \frac{2x}{3} +6-\color \red{\frac{2x}{3}} \\ 5x+\frac{x}{3}=6\\ 16x/3=6\\ 16x=18 \\ x=18/16\]

OpenStudy (anonymous):